if helen can ride 15km on her bike in the same time it takes her to walk 6km. If her rate riding is 6km/h faster than her rate walking how fast does she walk?
Answer by mananth(16946)

walking speed = x km/h
riding speed = x+6 km/h
distance riding = 15
and
distance walking = 6
time riding = 15/(x+6)
time walking = 6/x
both are same
6/x = 15/(x+6)
cross multiply
6(x+6)=15x
6x+36=15x
9x=36
/9
x=4
walking speed = 4km/h
riding speed = 6+4 = 10 km/h
m.ananth@hotmail.ca